OpenStudy (anonymous):

Which statement is true about the product square root of 2(3square root of 2 + square root of 18)? It is rational and equal to 12. It is rational and equal to 18. It is irrational and equal to 6 + 2square root of 18. It is irrational and equal to 3 +square root of 18.

jimthompson5910 (jim_thompson5910):

This expression? \[\Large \sqrt{2}*\left(3\sqrt{2}+\sqrt{18}\right)\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{2}*\left(3\sqrt{2}+\sqrt{18}\right)\] \[\Large \color{red}{\sqrt{2}*}\left(3\sqrt{2}+\sqrt{18}\right)\] \[\Large \color{red}{\sqrt{2}*}3\sqrt{2}+\color{red}{\sqrt{2}*}\sqrt{18}\] \[\Large 3*\sqrt{2}*\sqrt{2}+\sqrt{2}*\sqrt{18}\] \[\Large 3*\sqrt{2*2}+\sqrt{2*18}\] I'll let you finish up

OpenStudy (anonymous):

what am i supposed to do now? I'm kinda confused

jimthompson5910 (jim_thompson5910):

what is 2*2 equal to? what is 2*18 equal to?

OpenStudy (anonymous):

ohhh i understand thank you

jimthompson5910 (jim_thompson5910):

I'm glad it makes sense now